# Galley: Modern Query Optimization for Sparse Tensor Programs

### v3

@@ Abstract.
The tensor programming abstraction has become a foundational paradigm for modern computing. This framework allows users to write high performance programs for bulk computation via a high-level imperative interface. Recent work has extended this paradigm to sparse tensors (i.e. tensors where most entries are not explicitly represented) with the use of sparse tensor compilers. These systems excel at producing efficient code for computation over sparse tensors, which may be stored in a wide variety of formats. However, they require the user to manually choose the order of operations and the data formats at every step. Unfortunately, these decisions are both highly impactful and complicated, requiring significant effort to manually optimize. In this work, we present Galley, a system for declarative sparse tensor programming. Galley performs cost-based optimization to lower these programs to a logical plan then to a physical plan. It then leverages sparse tensor compilers to execute the physical plan efficiently. We show that Galley achieves high performance on a wide variety of problems including machine learning algorithms, subgraph counting, and iterative graph algorithms.
@@ 1. Introduction
In recent years, the tensor programming model has become ubiquitous across different fields of computation. Popularized by its use in deep learning, it has been applied to problems such as relational query processing (He et al., 2022; Asada et al., 2022; Koutsoukos et al., 2021), data cleaning (Sun et al., 2023), graph algorithms (Szárnyas et al., 2021), and scientific computing (Wang et al., 2022; Stonebraker et al., 2013; Modi et al., 2021) among others. This model promises a high-level imperative abstraction which is highly efficient as long as the problem can be posed in terms of tensor (i.e. array) operations. By doing so, it insulates the user from the many low level concerns that are crucial to achieving good performance. This has allowed experts in fields like data science and machine learning to take advantage of a wide range of hardware infrastructure without having to become experts in high performance computing as well.
Traditional tensor processing frameworks are built on collections of hand-optimized functions, called kernels, which each compute an operation over dense tensors (Abadi et al., 2016; Paszke et al., 2019; Harris et al., 2020; Anderson et al., 1990). The operation can be simple like a matrix-vector multiplication, or it can be a fusion of multiple semantic operations, like $A(B+C)$. However, most data is fundamentally sparse (i.e. most entries are a fill value like 00) such as graph data, one-hot encodings, relational data, meshes in physical simulations, sparse neural networks, and others. Even materializing these datasets as dense arrays can be prohibitively costly, so it’s crucial to perform the computation over its compressed, sparse format (Kjolstad et al., 2017).
To support sparse data, each framework needs to offer many distinct implementations of each operation, one for each combination of input formats. Each input can either be stored densely or in one of many sparse formats, e.g. CSR, CSC, COO. This has lead to an explosion of required kernels as the number of operations and formats both continue to grow. Understandably, these frameworks have been unable to keep up with this implementation effort, resulting in spotty coverage for operations over sparse data (Ivanov et al., 2023).
@@
Consider logistic regression over $n$ data points and $d$ features where $X\in\mathbb{R}^{n\times d}$ is the feature matrix and $\theta\in\mathbb{R}^{d}$ is the parameter vector. Inference is defined as
@@
where $\sigma$ is the sigmoid function. To compute $Y_{i}=\sum_{j}X_{ij}\theta_{j}$ in the dense case, one could use the ’matmul’ function from the Numpy library. On these inputs, this function specializes to an efficient, hand-coded implementation of dense matrix-vector multiplication. However, if any combination of $X$, $\theta$, and $Y$ can be sparse, Numpy needs to provide eight distinct implementations of matrix-vector multiplication.
To address this issue, the scientific computing community has adopted sparse tensor compilers (STCs) (Kjolstad et al., 2017; Bik et al., 2022; Ahrens et al., 2024; Shaikhha et al., 2022; Hu et al., 2019). These compilers take in a high level imperative tensor program and, separately, a description of the input tensors’ formats [^1] Some systems separate declarative and imperative concerns with a scheduling language. However, the user still controls both aspects.. They automatically produce an efficient kernel implementation in low level code (LLVM, MLIR, etc.). Thus, STCs offer a form of data independence, by allowing the user to separately specify the algorithm and the data layout. For example, Fig. 1 shows a kernel definition for Ex. 1.1 written in Finch, an STC language (Ahrens et al., 2024). The Dense/Sparse input formats are specified in line 1 and are separate from the imperative code in lines 5-15.
However, STCs require users to make a series of complex decisions to produce an efficient implementation. First, the user needs to break their program into a sequence of aggregations. Suppose that the feature matrix in Ex. 1.1 is the result of a matrix multiplication, $X=A\cdot B$, and Eq. (1) becomes $Y_{i}=\sigma(\sum_{jk}A_{il}B_{lj}\theta_{j})$. The user must choose to sum over $j$ or $l$ first, resulting in vastly different runtimes (Ahrens et al., 2022; Zhang et al., 2021). Then, for each aggregate, the user needs to choose the output format, loop order, and iteration algorithm. Consider Fig 1. The user must choose the output format for the intermediate R (line 3). In this case, she chose a Dense format rather than a Sparse format which would be 10$\approx$ 10$\times$ 10 × slower. Then, the user chooses the loop order (lines 5-6). She chose $i$-then-$j$ which is asymptotically faster than $j$-then-$i$ because each out-of-order access to $X$ requires a full scan of the tensor. Lastly, she picks a merge algorithm for each loop that describes how to iterate through the non-zero indices (line 8). Here, $X$ is iterated through and each non-zero $j$ is looked up in $\theta$. If she chose to iterate through $\theta$, each inner loop would scan the entire vector. While STCs allow users to separate the algorithm from the data formats, the user is still responsible for optimizing her program to achieve good performance.
@@ 15. END
Figure 1. Logistic Regression Impl. in Simplified Sparse Tensor Compiler Language.
@@
In this paper, we propose Galley, a system for declarative sparse tensor programming. Users write tensor programs in a declarative language based on the Einsum notation, similar to Equation (1), and Galley automatically produces an optimized STC implementation. First, it restructures the program into a sequence of aggregation steps, minimizing the total computation and materialization costs (Sec. 5). Then, Galley optimizes each step by selecting the optimal formats for all intermediate tensors, the loop order, and the access method for each index access (Sec. 6). This is all guided by a system for estimating sparsity via statistics on the input tensors (Sec. 7). Galley builds on fundamental principles from cost-based query optimization while developing new techniques that are specific to producing optimized code for sparse tensor compilers(Khamis et al., 2016).
@@
In the first phase, Galley rewrites the input program into a sequence of aggregation steps by adapting the variable elimination framework. This reduces a complex rewriting problem to the problem of finding an optimal elimination order. However, this approach has previously only been defined for sum-product queries. To handle arbitrary sparse tensor algebra programs, we extend it to non-distributive functions (e.g. $\sum_{i}max(A_{ij},B_{j})$), disjunctive functions (e.g. $\sum_{i}A_{i}+B_{i}=\sum_{i}A_{i}+\sum_{i}B_{i}$), and internal aggregates (e.g. $\sum_{j}A_{j}\sqrt{\sum_{k}B_{jk}}$). In Sec. 9, we show that these optimizations produce up to a $\mathbf{100\times}$ speedup for machine learning algorithms over composite feature matrices. After this phase, our program has been converted to a sequence of aggregates over pointwise expressions, i.e. the Logical Plan dialect in Fig. 4.
@@
Next, Galley’s physical optimizer transforms each aggregate into an efficient STC kernel by choosing the loop order, output format, and merge algorithm. Because sparse outer loops avoid inner iterations, a good loop ordering can significantly improve performance. Galley finds the optimal loop order with a novel combination of branch-and-bound and dynamic programming techniques. Following the modern tensor format abstraction (i.e. the fibertree abstraction), Galley then chooses a format for each dimension of the output by examining both the sparsity and write pattern (random vs sequential)(Sze et al., 2017; Chou et al., 2018; Bik et al., 2022). Lastly, for each loop index, we select one input to iterate over and perform lookups on the others.
@@
All of Galley’s optimizations are guided by a framework for estimating the sparsity of intermediate results from statistics on the inputs. The core of this is a minimal interface for composing statistics about inputs to infer statistics about outputs. This makes incorporating a new kind of sparsity statistic as simple as implementing 5 functions to; 1) collect statistics on inputs 2-3) handle conjunctive and disjunctive point-wise operations 4) handle aggregates and 5) produce sparsity estimates. We demonstrate this by implementing a naive estimator based on uniformity assumptions and a worst-case estimator based on cardinality bounds.
@@
In Example 1.1, we assumed that $X$ is given directly. However, feature matrices are generally assembled from more basic inputs. Suppose $S_{ipc}\in\mathbb{B}^{n_{i}\times n_{p}\times n_{c}}$ is a boolean sparse tensor where each non-zero entry is a sale, $i$, of product, $p$, to customer, $c$. $P_{pj}\in\mathbb{R}^{n_{p}\times d}$ and $C_{cj}\in\mathbb{R}^{n_{c}\times d}$ are smaller matrices holding numeric features about products and customers. The feature dimension, indexed by $j$, represents both product and customer features, but each feature comes from either $P_{pj}$ or $C_{cj}$. So, each column is non-zero in either $P$ or $C$, and they are concatenated by addition. $X$ is now defined as $X_{ij}=\sum_{pc}S_{ipc}(P_{pj}+C_{cj})$. Galley’s logical optimizer can take advantage of $X$’s structure by pushing $\theta$ down into its definition,
$Y_{i}=\sigma(\sum_{jpc}S_{ipc}P_{pj}\theta_{j}+\sum_{jpc}S_{ipc}C_{cj}\theta_{j})$
We can then push down the summations,
(2) $Y_{i}=\sigma\left(\sum_{pc}\left(S_{ipc}\sum_{j}\left(P_{pj}\theta_{j}\right)\right)+\sum_{pc}\left(S_{ipc}\sum_{j}\left(C_{cj}\theta_{j}\right)\right)\right)$
By doing this, Galley only materializes vector intermediates rather than the full feature matrix. On similar examples, Sec. 9 shows that this can result in up to 100×100$\times$100 × faster execution.
After logical optimization, Galley’s physical optimizer produces a single STC kernel for each aggregate. For $\sum_{j}P_{pj}\theta_{j}$, it would produce,
@@
Contributions We claim the following contributions in this work,
  • We present Galley, a system for declarative sparse tensor programming (Sec.4). Galley automatically translates high-level sparse tensor programs to efficient imperative programs targeting a sparse tensor compiler.
  • Galley supports a highly expressive language for sparse tensor algebra with arbitrary algebraic operators, aggregates within expressions, and multiple outputs.
  • Galley is the first system to perform cost-based logical optimization, redistributing aggregates and reorganizing the pointwise statements in the program. (Sec.5)
  • Galley is the first system to perform cost-based physical optimization to determine loop orders, tensor formats, and merge algorithms for each dimension (Sec.6)
  • We propose a minimal interface for sparsity estimation and demonstrate it by implementing two estimators (Sec.7)
  • We evaluate Galley on several workloads and show that it is up to 100x faster than hand-optimized kernels for mixed dense-sparse workloads and up to 100x faster than a state of the art database for highly sparse workloads.
Figure 2. Fibertree Format Abstraction
@@ 2. Related Work
Our work differs from other proposals on cost-based optimization for tensor processing due to its targeting of STCs and its expressive input language. SystemDS, formerly SystemML, is a system for end-to-end machine learning over matrices and tabular data (Boehm et al., 2016, 2020; Baunsgaard and Boehm, 2023). It primarily takes in linear algebra (LA) programs and targets a combination of LA libraries and distributed computing via Spark. Later work, SPORES, extended its logical optimizer to leverage relational algebra when optimizing sum-product expressions(Wang et al., 2020). Their core insight was that LA rewrites, which always match and produce 0-2D expressions, are not sufficient. Optimal rewrites need to pass through higher order intermediate expressions. Other work translated sum-product expressions to SQL to leverage highly efficient database execution engines(Blacher et al., 2023). These systems can perform well for highly sparse inputs but struggle on mixed dense-sparse workloads. Tensor relational algebra proposes a relational layer on top of dense tensor algebra which provides a strong foundation for automatically optimizing distributed dense tensor computations (Yuan et al., 2021). In the compilers community, there have been attempts to automatically optimize sparse tensor kernels based on asymptotic performance analyses(Ahrens et al., 2022; Dias et al., 2023). These systems each target a different execution context and focus on different aspects of optimization. Galley expands on this line of work by targeting a new execution engine, proposing novel optimization techniques, and handling a wider range of tensor programs.

### v4

@@ Abstract.
The tensor programming abstraction has become a foundational paradigm for modern computing. This framework allows users to write high performance programs for bulk computation via a high-level imperative interface. Recent work has extended this paradigm to sparse tensors (i.e., tensors where most entries are not explicitly represented) with the use of sparse tensor compilers. With these systems, users define the semantics of the program and the algorithmic decisions in a concise language that can be compiled to efficient low-level code. However, these systems place users in the role of performance engineers, requiring them to make complex decisions about program structure and memory layouts to write programs that run efficiently.
This work presents Galley, a system for declarative tensor programming that allows users to write efficient tensor programs without making complex algorithmic decisions. Galley is the first system to perform cost based lowering of sparse tensor algebra to the imperative language of sparse tensor compilers, and the first to optimize arbitrary operators beyond $\sum$ and *. It does this with a two-level, cost-based program optimizer. At the logical level, it decomposes the input program into a sequence of aggregation steps through a novel extension of the FAQ framework. At the physical level, Galley optimizes and converts each aggregation step to a concrete program, which is compiled and executed with a sparse tensor compiler. We show that Galley produces programs that are 1300×1-300$\times$1 - 300 × faster than competing methods for machine learning over joins and 520×5-20$\times$5 - 20 × faster than a state-of-the-art relational database for subgraph counting workloads. Finally, we show that Galley introduces a minimal optimization overhead across all workloads.
@@ 1. Introduction
In recent years, the tensor programming model has become ubiquitous across different fields of computation. Popularized by its use in deep learning, it has been applied to problems such as relational query processing (He et al., 2022; Asada et al., 2022; Koutsoukos et al., 2021), data cleaning (Sun et al., 2023), graph algorithms (Szárnyas et al., 2021), and scientific computing (Wang et al., 2022; Stonebraker et al., 2013; Modi et al., 2021), among others. This model promises a high-level imperative abstraction that is highly efficient as long as the problem can be posed in terms of tensor (i.e., array) operations. It thereby insulates the user from the many low-level concerns that are crucial to a program’s efficient performance. This, in turn, has allowed experts in fields like data science and machine learning (ML) to leverage a wide range of hardware infrastructures without having to become expert in high-performance computing.
Traditional tensor processing frameworks are built on collections of hand-optimized functions, called kernels, which each compute an operation over dense tensors (Abadi et al., 2016; Paszke et al., 2019; Harris et al., 2020; Anderson et al., 1990). The operation can be simple, like a matrix-vector multiplication, or it can be a fusion of multiple semantic operations, like $A(B+C)$. However, most data is fundamentally sparse (i.e., most entries are a fill value like 00), including graph data, one-hot encodings, relational data, 3D physics meshes, sparse neural networks, and others. Even materializing these datasets as dense arrays can be prohibitively costly, so it is crucial to perform the computation over its compressed, sparse format (Kjolstad et al., 2017).
To support sparse data, each framework offers many distinct implementations of each operation, one for each combination of input formats. Each input can either be stored densely or in one of many sparse formats, e.g., CSR, CSC, COO. This has lead to an explosion of required kernels as the number of operations and formats continue to multiply. Understandably, these frameworks have been unable to keep pace with this implementation effort, resulting in spotty coverage for operations over sparse data (Ivanov et al., 2023).
@@
Consider logistic regression over $n$ data points and $d$ features, where $X\in\mathbb{R}^{n\times d}$ is the feature matrix and $\theta\in\mathbb{R}^{d}$ is the parameter vector. Inference is defined as
@@
where $\sigma$ is the sigmoid function. To compute $Y_{i}=\sum_{j}X_{ij}\theta_{j}$ in the dense case, one could use the ’matmul’ function from the Numpy library. On these inputs, this function specializes to an efficient, hand-coded implementation of dense matrix-vector multiplication. However, if any combination of $X$, $\theta$, and $Y$ could be sparse, Numpy needs to provide eight distinct implementations of matrix-vector multiplication.
To address this issue, the scientific computing community has adopted sparse tensor compilers (STCs) (Kjolstad et al., 2017; Bik et al., 2022; Ahrens et al., 2024; Shaikhha et al., 2022; Hu et al., 2019). These compilers take as input a high-level imperative tensor program and a separate description of the input tensors’ formats. [^1] Some systems separate declarative and imperative concerns with a scheduling language. However, the user still controls both aspects. They automatically produce an efficient kernel implementation in low-level code (LLVM, MLIR, etc.). Thus, STCs offer a form of data independence by letting users separately specify the algorithm and the data layout. For example, Fig. 1 shows a kernel definition for Ex. 1.1 written in Finch, an STC language (Ahrens et al., 2024). The Dense/Sparse input formats are specified in line 1 and are separate from the imperative code in lines 5-15.
However, STCs require users to make a series of complex decisions to produce an efficient implementation. First, users must break their program into a sequence of aggregations. Suppose that the feature matrix in Ex. 1.1 is the result of a matrix multiplication $X=A\cdot B$, so Eq. (1) becomes $P_{i}=\sigma(\sum_{jk}A_{il}B_{lj}\theta_{j})$. Users must choose to sum over $j$ or $l$ first, resulting in vastly different runtimes (Ahrens et al., 2022; Zhang et al., 2021). Then, for each aggregate, they must choose the output format, loop order, and iteration algorithm.
Consider Fig 1. Here, the user must choose the output format for the intermediate R (line 3). In this case, she chose a Dense rather than a Sparse format, which would be 10$\approx$ 10$\times$ 10 × slower. Then, the user chooses the loop order (lines 5-6). In this case, she chose $i$-then-$j$, which is asymptotically faster than $j$-then-$i$ because each out-of-order access to $X$ requires a full scan of the tensor. Finally, the user picks a merge algorithm for each loop that describes how to iterate through the non-zero indices (line 8). Here, $X$ is iterated through, and each non-zero $j$ is looked up in $\theta$. If she chose to iterate through $\theta$, each inner loop would scan the entire vector. Thus, while STCs let them separate the algorithm from the data formats, users are still responsible for optimizing their program to achieve efficient performance.
@@ 15. END
Figure 1. Logistic regression implemented in the language of a sparse tensor compiler.
@@
In this paper, we propose Galley, a system for declarative sparse tensor programming. Galley makes algorithmic decisions on the users’ behalf, freeing them to focus on the high-level semantics of their program without sacrificing computational efficiency. It accepts input programs written in a declarative language based on Einsum notation, similar to Equation (1), and automatically produces an optimized STC implementation. To do so, it first restructures the program into a sequence of aggregation steps, minimizing total computation and materialization costs (Sec. 5). It then optimizes each step by selecting the loop order, the optimal formats for all intermediate tensors, and the intersection algorithm for each loop (Sec. 6). These decisions are all guided by a system for estimating sparsity via statistics on the input tensors (Sec. 7). Galley builds on fundamental principles from cost-based query optimization while developing new techniques that are specific to producing optimized code for sparse tensor compilers(Khamis et al., 2016).
@@
In the first phase of program optimization, Galley’s logical optimizer rewrites the input program into a sequence of aggregation steps using our new extension of the variable elimination (VE) framework. By adopting this framework, we reduce a complex rewriting problem to one of finding the optimal order to aggregate indices. However, VE has previously been defined only for sum-product expressions, while Galley lets users write arbitrary sparse tensor algebra programs. These programs include non-distributive functions (e.g., $\sum_{i}max(A_{ij},B_{j})$), non-commutative aggregates (e.g., $\sum_{j}\max_{k}B_{jk}$), and other constructs that break the typical assumptions of VE.
Fortunately, we show that this complexity can be neatly managed by reasoning over the annotated expression tree, which lets Galley benefit from algebraic properties when they exist and respect them when they do not. In Sec. 8, we show that these logical optimizations produce up to a $\mathbf{100\times}$ speedup for ML algorithms over composite feature matrices. At the end of this phase, our program has been converted to a sequence of aggregates over pointwise expressions, i.e., the Logical Plan dialect in Fig. 4.
@@
Next, Galley’s physical optimizer generates an efficient STC kernel for each aggregate by choosing the loop order, output format, and merge algorithm. Because sparse outer loops avoid inner iterations, an efficient loop ordering can significantly improve performance. Galley finds the optimal loop order using a novel combination of branch-and-bound and dynamic programming techniques. Following the modern tensor format abstraction (i.e., the fibertree abstraction), Galley then chooses a format for each dimension of the output by examining both the sparsity and write pattern (random vs sequential)(Sze et al., 2017; Chou et al., 2018; Bik et al., 2022). Finally, for each loop index, it selects one input to iterate over and performs lookups on the others.
@@
Galley’s optimizations are guided by a sparsity estimation framework. At its core is a minimal interface for composing statistics. To incorporate a new sparsity estimator in Galley, we only need to implement 5 functions, those for: (1) collecting statistics on inputs, (2-3) handling conjunctive (i.e., *) and disjunctive (i.e., $+$) point-wise operations, (4) handling aggregates, and (5) producing sparsity estimates. We demonstrate this by implementing a uniform sparsity estimator and a worst-case estimator based on cardinality bounds.
In Example 1.1, we assumed that $X$ is given directly. However, feature matrices are generally assembled from more basic inputs. As we show in the following example, Galley can take advantage of this kind of program structure to produce superior performance.
@@
Assume $S_{ipc}\in\mathbb{B}^{n_{i}\times n_{p}\times n_{c}}$ is a boolean sparse tensor, where each non-zero entry is a sale, $i$, of product, $p$, to customer, $c$. $P_{pj}\in\mathbb{R}^{n_{p}\times d}$ and $C_{cj}\in\mathbb{R}^{n_{c}\times d}$ are smaller matrices holding numeric features about products and customers. The feature dimension, indexed by $j$, represents both product and customer features, but each feature comes from either $P_{pj}$ or $C_{cj}$. So, each column is non-zero in either $P$ or $C$, and they are concatenated by addition. $X$ is now defined as $X_{ij}=\sum_{pc}S_{ipc}(P_{pj}+C_{cj})$. Galley’s logical optimizer can take advantage of $X$’s structure by pushing $\theta$ down into its definition, as follows.
$Y_{i}=\sigma(\sum_{jpc}S_{ipc}(P_{pj}\theta_{j}+C_{cj}\theta_{j}))$
We can then push down the summation over $j$, as follows,
(2) $Y_{i}=\sigma(\sum_{pc}S_{ipc}(\sum_{j}P_{pj}\theta_{j}+\sum_{j}C_{cj}\theta_{j}))$
In this way, Galley materializes only vector intermediates, not the full feature matrix. On similar examples, Sec. 8 shows that such changes can yield up to 300×300$\times$300 × faster execution.
After logical optimization, Galley’s physical optimizer produces a single STC kernel for each aggregate. For $\sum_{j}P_{pj}\theta_{j}$, it would produce
@@
Contributions We claim the following contributions:
  • We present Galley, a system for declarative sparse tensor programming (Sec.4). Galley is the first system to perform cost based lowering of sparse tensor algebra to the imperative language of sparse tensor compilers, and the first to optimize arbitrary operators beyond $\sum$ and *.
  • Galley supports a highly expressive language for sparse tensor algebra with arbitrary algebraic operators, aggregates within expressions, and multiple outputs (Sec.4).
  • Galley performs cost-based logical optimization with a novel extension of the variable elimination framework to handle arbitrary aggregations and pointwise operators (Sec.5). Galley performs cost-based physical optimization to determine loop orders, tensor formats, and merge algorithms for each dimension (Sec.6).
  • We propose a minimal interface for sparsity estimation to guide optimizations and demonstrate it by implementing two estimators (Sec.7).
  • We evaluate Galley and show that it is 1-300x faster than hand-optimized kernels for mixed dense-sparse workloads and 5-20x faster than a state-of-the-art database for highly sparse workloads based on several evaluation workloads (Sec.8).
@@ 2. Related Work
Galley differs from other work on cost-based optimization for tensor processing due to its targeting of STCs and its expressive input language. SystemDS, formerly SystemML, focuses on end-to-end ML over matrices and tabular data (Boehm et al., 2016, 2020; Baunsgaard and Boehm, 2023); it takes as input linear algebra (LA) programs and targets a combination of LA libraries and distributed computing via Spark. Later work, SPORES, extended its logical optimizer to leverage relational algebra when optimizing sum-product expressions(Wang et al., 2020); their core insight was that LA rewrites, which always match and produce 0-2D expressions, are not sufficient and that optimal rewrites must pass through higher order intermediate expressions. Other related work translated sum-product expressions to SQL to leverage highly efficient database execution engines (Blacher et al., 2023). These systems can perform well for highly sparse inputs but struggle on mixed dense-sparse workloads. Tensor relational algebra proposes a relational layer on top of dense tensor algebra that provides a strong foundation for automatically optimizing distributed dense tensor computations (Yuan et al., 2021; Bourgeois et al., 2024). The compiler community has made attempts to automatically optimize sparse tensor sum-product kernels based on asymptotic performance analyses(Ahrens et al., 2022; Dias et al., 2023). These systems each target a different execution context and focus on different aspects of optimization. Galley expands on this line of work by targeting a new execution engine, proposing novel optimization techniques, and handling a wider range of tensor programs.